OpenStudy (anonymous):

IM GIVING MEDALS TO THE BEST HELPER Point P is located at (−2, 7) and point R is located at (1, 0). Find the y value for the point Q that is located two over three the distance from point P to point R. 4.9 4.7 2.5 2.3

OpenStudy (anonymous):

HELLLP

OpenStudy (anonymous):

Try the point Q=(-4/5, 21/5)

OpenStudy (anonymous):

I will explain to you how you get it.

OpenStudy (anonymous):

Any point between P and R can be written as \[ g(t)= P + t(R-P)= (-2,7) +t((1,0)-(-2,7))= \\(-2 + 3 t, 7-7t) \quad 0 \le t\le 1 \]

OpenStudy (anonymous):

For example g(0)=P; g(1)=R, g(1/2) is the midpoint of the segment PR

OpenStudy (anonymous):

to have the point Q so that \( \frac{PQ}{QR} =2/3 \), we need to divide Pr into 5 equal parts, and we should find PQ=2/5 and QR=3/5. Hence Q=g[2/5)=(-4/5,21/4)

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

YW